OpenStudy (anonymous):

If a,b,c are in A.P. prove that a^2+c^2+4ac=2(ab+bc+ca) now as a,b,c are in AP b-a=c-b or 2b = a+c, but can't proceed further..

OpenStudy (theeric):

I'm not sure if I can help... What is the AP set?

OpenStudy (anonymous):

@terenzreignz : could you please help me with this ?

terenzreignz (terenzreignz):

Interesting how they contain squares... have you tried squaring both sides of 2b = a+c ?

OpenStudy (ankit042):

add b^2 on both the side and move the right side to left..you can see (a-b+c)^2=b^2

terenzreignz (terenzreignz):

This is going to take some creativity...which I lack :D hang on

OpenStudy (anonymous):

so @ankit042 both sides will equate to b^2

OpenStudy (ankit042):

Yeah you have to show L.H.S = R.H.S

OpenStudy (anonymous):

thanks @ankit042